# FLARE: Faithful Logic-Aided Reasoning and Exploration

### v4

@@ Abstract
Modern Question Answering (QA) and Reasoning approaches based on Large Language Models (LLMs) commonly use prompting techniques, such as Chain-of-Thought (CoT), assuming the resulting generation will have a more granular exploration and reasoning over the question space and scope. However, such methods struggle with generating outputs that are faithful to the intermediate chain of reasoning produced by the model. On the other end of the spectrum, neuro-symbolic methods such as Faithful CoT (F-CoT) and Logic-LM propose to combine LLMs with external symbolic solvers. While such approaches boast a high degree of faithfulness, they usually require a model trained for code generation and struggle with tasks that are ambiguous or hard to formalise strictly. We introduce Faithful Logic-Aided Reasoning and Exploration (FLARE), a novel interpretable approach for traversing the problem space using task decompositions. We use the LLM to plan a solution, formalise the query into facts and predicates, which form the problem space, using a logic programming code and simulate that code execution using an exhaustive multi-hop search over the defined space. Our method allows us to compute the faithfulness of the reasoning process w.r.t. the generated code and explicitly trace the steps of the multi-hop search without relying on external solvers. Our methods achieve SOTA results on $\mathbf{7}$ out of $\mathbf{9}$ diverse reasoning benchmarks. We also show that model faithfulness positively correlates with overall performance and further demonstrate that FLARE allows pinpointing the decisive factors sufficient for and leading to the correct answer with optimal reasoning during the multi-hop search. Our findings reveal that successful traces exhibit, on average, a $18.1\%$ increase in unique emergent facts, a $8.6\%$ higher overlap between code-defined and execution-trace relations, and a $3.6\%$ reduction in unused code relations.
@@ 1 Introduction
Complex Reasoning in natural Question Answering (QA) tasks assumes the capability to explore the problem space of the designated query with a formalised set of facts, relations, commonsense knowledge and logical implications. In line with this, LLMs have been enhanced with CoT (Wei et al., 2022) prompting, which supplements the QA process by generating intermediate reasoning chains given a set of in-context examples (Brown et al., 2020a), as shown in Figure 1. This allowed for advancement in commonsense (Madaan et al., 2022), symbolic (Wang et al., 2022; Sprague et al., 2024) and mathematical (Jie et al., 2023) reasoning. Although CoT allows for a problem exploration in natural language steps, such an approach has been shown to cause performance degradation for reasoning tasks involving multi-step planning (Valmeekam et al., 2022; Suzgun et al., 2023), problem exploration (Yao et al., 2022), and arithmetic tasks (Hendrycks et al., 2021b; Madaan & Yazdanbakhsh, 2022b). These discrepancies arise as CoT suffers from a limited ability to decompose, search, verify and backtrack using intermediate rationale chains (Yao et al., 2022), cascading hallucinations and errors (Ling et al., 2023) and that natural language might not be an optimal representation for describing the reasoning process (Li et al., 2024). Simultaneously, LLM output has been shown to be unfaithful and inconsistent w.r.t. the intermediate CoT rationale (Jacovi et al., 2024; Lanham et al., 2023b; Turpin et al., 2023).
@@
To overcome these problems, we propose Faithful Logic-Aided Reasoning and Exploration (FLARE), an interpretable method that allows for planning, fuzzy reasoning, and traversing the problem space with backtracking, exact task decomposition, and measuring faithfulness. In FLARE, given a natural language query, we prompt an LLM to sequentially generate a plan that includes an analysis and the logical steps necessary for formalising and answering the question, a logic programming (Wielemaker et al., 2012) code that allows formalising the query into a set of facts, relations and their composition forming the space for exploring that query and the search, which is an LLM-generated code execution simulation. An illustration of FLARE can be seen in Figure 1. In our framework, the generated code must not be consistently executable by an external solver, allowing for the soft-formalisation of natural language. Although we see that even generalist LLMs are able to produce executable code in $\geq 50\%$ of cases. FLARE allows us to measure the faithfulness of the outcome w.r.t. the simulated code execution by directly comparing the search paths produced by the external solver to that LLM generation. This comparison also allows for pinpointing model hallucinations and inconsistencies. We systematically study the effectiveness of our method using $4$ general-purpose LLMs of varying scales across $9$ diverse QA and $3$ logical inference benchmarks, covering Math World Problems, Multi-hop QA, Relation inference, deductive and analytical reasoning and show that our method achieves state-of-the-art results in $7$ out of $9$ QA datasets and $2$ out of $3$ logic datasets in comparison to CoT, F-CoT and Logic-LM. We also show that the method is competitive for models tuned for coding, with an average overall increase of $16\%$ over F-Cot and $9\%$ over CoT. Our findings show that model accuracy strongly correlates with the faithfulness of the reasoning process towards search traces from the simulated code execution. We also provide ablations showing that the model can interpretably pinpoint hallucinations, underutilized knowledge, and the limitations of the search over the problem space. Our key contributions are the following:
  • We introduce FLARE a novel paradigm for logic-aided and interpretable formalisation and search over the problem space in QA and logic reasoning tasks.
  • We perform a systematic evaluation across $9$ QA and $3$ logical inference benchmarks and $4$ models of varying scales, showing the advantages of using FLARE for QA in a few-shot setup over prior approaches.
  • The modularity of FLARE allows defining a simple ingrained method for measuring model faithfulness, which is further shown to be strongly correlated with performance.
  • We further show that using FLARE allows us to interpretably and rigorously detect hallucinations along with sub-optimal and inconsistent reasoning patterns.
@@ 2 Related Work
Few-shot prompting (Brown et al., 2020b) has been shown to be an effective approach for increasing the reasoning capabilities of LLMs in natural language generation (Gehrmann et al., 2021; Reif et al., 2022; Sanh et al., 2022). LLM reasoning can be further enhanced with prompting techniques such as CoT (Wei et al., 2022), which attempts to segment reasoning into explicitly written intermediate steps. Concurrent work has also proposed that models “think step by step” (Kojima et al., 2022), or divide the problem into subtasks before the solution (Zhou et al., 2023, Least-to-Most). These approaches have been shown to suffer from arithmetic inaccuracies (Lewkowycz et al., 2022; Hendrycks et al., 2021a) and reasoning inconsistencies (Madaan & Yazdanbakhsh, 2022a). Further attempts have been made to add a planning stage before reasoning by dividing the process into recursive plan formulation and execution steps (Yao et al., 2023b; Wang et al., 2023a). The plan generation step in FLARE is a hybrid technique inspired by these methods but focused on generating a natural language strategy for formalising the query into code.
@@
Several lines of work propose using techniques to expand the reasoning paths over the problem space. Self-consistency decoding (Wang et al., 2023b) is an approach used to sample many natural language reasoning paths and take a majority vote for an answer. Another popular approach is Tree-of-Thoughts (ToT; Yao et al., 2023a), which proposes to explore the query with reasoning similar to a tree traversal, where each state is created and evaluated using an LLM. Similar techniques try to adapt symbolic search approaches akin to DFS, BFS (Besta et al., 2024), A (Lehnert et al., 2024) or other combinations (Gandhi et al., 2024) with direct tuning (Lehnert et al., 2024), imitation training (Yang et al., 2022) or few-shot prompting (Zhang et al., 2024). It must be noted that all of these techniques have only been tested in constrained mathematical puzzle-solving and algorithmic domains like the 24 Game (Yang et al., 2022), Countdown (Wikipedia, 2024), Sorting (Besta et al., 2024), maze solving (Yang et al., 2022), Sokoban (Lehnert et al., 2024), and others. Although the search component of FLARE has some similarities to these techniques, we argue that our method allows for generalistic reasoning with interpretable multi-hop search through simulated code execution.
@@
Another line of research has tried formalising natural language queries into code (Gao et al., 2023; Li et al., 2024) or pseudo-code (Chae et al., 2024; Gandhi et al., 2024). This allows the translation of the query into a strict structure and offloads the reasoning and search components to deterministic solvers like Python Chen et al. (2023), PDDL Lyu et al. (2023); Liu et al. (2023), DataLog Lyu et al. (2023) and others. While models are capable of synthesising programs (Austin et al., 2021; Nijkamp et al., 2023) and benefit from the use of code in numerical and algorithmic reasoning settings (Chen et al., 2023; Gao et al., 2023), the usage of code for general QA has not been rigorously explored. The reasons are that formalisation from natural language into a strict and executable code is challenging (Wu et al., 2022), following the exact syntactic constraints of the programming language not abundantly used during pre-training is onerous (Liu et al., 2024) and can require models explicitly tuned for coding (Chen et al., 2021). Using an external solver for reasoning also limits the capability for soft reasoning in commonsense knowledge and implications. Although we formalise the natural language query into a logic programming Prolog program during the code generation part of FLARE, we do not explicitly require the code to be executable and do not use external solvers during inference. This allows for the further use of the LLM for soft-reasoning to simulate code execution in a logic-based problem space traversal similar to Prolog while circumventing the need for code tuning a generalist model.
@@
An explanation is considered faithful if it explicitly and accurately describes the reasoning process of the model during inference (Gilpin et al., 2018; Jacovi & Goldberg, 2020). In the context of prompting techniques such as CoT, we are interested in the faithfulness of the intermediate reasoning chains towards the final output. Faithful intermediate reasoning chains should not just look plausible (Herman, 2017) but have exact reflections of the problem exploration and reasoning used to arrive at the final answer. Natural language reasoning chains prevalent in CoT and similar methods are shown to be unfaithful, either masking the reasoning biases (Turpin et al., 2023) of the model or outright ignoring the intermediate reasoning (Lanham et al., 2023a). In FLARE, we introduce a method to seamlessly measure the faithfulness of the final outcome w.r.t. completed search.
@@ 3.1 LLM Simulated Search
For each query $\mathcal{Q}$, given an LLM $\mathcal{M}$, we initially use instructions $\mathcal{I}^{\mathcal{P}}$ to prompt it to generate a plan $\mathcal{P}$, which should be comprised of task explanation, analysis and a plan for further formalising the query. An example of this can be seen in the plan section in Figure 1. We use in-context few shot examples $\mathcal{E}_{\mathcal{P}}$ of such plan generations along with greedy decoding for obtaining the final plan.
$\mathcal{P}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{P}}_{i}\mid T^{\mathcal{P}}_{:i-1},\mathcal{E}_{\mathcal{P}},\mathcal{Q},\mathcal{I}^{\mathcal{P}})$ (1)
Where $\mathcal{P}_{i}\text{ and }T^{\mathcal{P}}_{i}$ is the $i$-th token in the generated plan $\mathcal{P}$ and $p_{\mathcal{M}}$ is the probability of the next token over the vocabulary obtained from model $\mathcal{M}$.
@@
After generating the plan, we use instructions $\mathcal{I}^{\mathcal{C}}$ to prompt the LLM $\mathcal{M}$ to generate a Prolog code $\mathcal{C}$, an example of which can be seen in Figure 1. We append executable code generation samples $\mathcal{C}_{\text{sample}}$ to the previous in-context examples $\mathcal{E}_{\mathcal{P}}$ and obtain few-shot code generation demonstrations $\mathcal{E}_{\mathcal{C}}=[\mathcal{E}_{\mathcal{P}};\mathcal{C}_{\text{sample}}]$
$\mathcal{C}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{C}}_{i}\mid T^{\mathcal{C}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}})$ (2)
$\mathcal{F}_{\text{\emph{code}}},\mathcal{R}_{\text{\emph{code}}},\mathcal{G}_{\text{\emph{code}}}=\text{EXTRACT}(\mathcal{C}_{i})$
Where $\mathcal{C}_{i}\text{ and }T^{\mathcal{C}}_{i}$ is the $i$-th token in the generated code $\mathcal{C}$.
@@
Benefits of Prolog
Prolog is a symbolic logic-programming engine (Bowen, 1979) designed for heuristic search over Horn Clauses (Chandra & Harel, 1985). As a declarative programming paradigm (Lloyd, 1994), the code is expressed as the logic of computation, expressed as a set of facts $\mathcal{F}$ and relations $\mathcal{R}$ defining the problem space, with the goal $\mathcal{G}$ being a first-order logic combination of them. Prolog employs depth-first search (DFS) (Bowen, 1979) for sub-goal decomposition and traversal of the problem space, satisfying $\mathcal{G}$ through a sequence of steps known as the trace. Each step either confirms/invalidates a sub-goal, expands the search tree, or retries failed sub-goals with new combinations. An example of such a search is shown in Figure 1. Prolog supports exhaustive search by exploring all paths that satisfy or fail the goal. This explicit segmentation of facts, relations, and search strategies simplifies query formalization. As a declarative language, Prolog enables segmentation using a simple regexp heuristic, referred to as EXTRACT in Equation 2 and Equation 3. Including exhaustive traces in-context allows an LLM to simulate sub-goal decomposition, backtracking, and intermediate goal invalidation, discussed further in the next paragraph.
@@
After generating the logic-programming code, we want to simulate program execution by generating a problem space traversal trace with our LLM $\mathcal{M}$. We use instructions $\mathcal{I}^{\mathcal{S}}$ and update our in-context samples by appending search traces $\mathcal{S}_{\text{sample}}$ constructed from Prolog execution of sample codes $\mathcal{C}_{\text{sample}}$, i.e. $\mathcal{E}_{\mathcal{S}}=[\mathcal{E}_{\mathcal{C}};\mathcal{S}_{\text{sample}}]$.
$\mathcal{S}_{i}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{S}}_{i}\mid T^{\mathcal{S}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}},\mathcal{C},\mathcal{I}^{\mathcal{S}})$ (3)
$\mathcal{A}_{\text{\emph{search}}},\mathcal{F}_{\text{\emph{search}}},\mathcal{R}_{\text{\emph{search}}}=\text{EXTRACT}(\mathcal{S}_{i})$
@@
$\mathcal{A}_{\text{\emph{Final}}}\sim\operatorname*{arg\,max}p_{\mathcal{M}}(T^{\mathcal{A}}_{i}\mid T^{\mathcal{A}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}},\mathcal{C},\mathcal{I}^{\mathcal{S}},\mathcal{S},\mathcal{I}^{\mathcal{A}})$ (4)
The prompts used for generating each part in FLARE can be seen in Appendix A along with a complete example in Table 8 and a pseudo-code in Algorithm 1.

### v5

@@ Abstract
Modern Question Answering (QA) and Reasoning approaches with Large Language Models (LLMs) commonly use Chain-of-Thought (CoT) prompting but struggle with generating outputs faithful to their intermediate reasoning chains. While neuro-symbolic methods like Faithful CoT (F-CoT) offer higher faithfulness through external solvers, they require code-specialized models and struggle with ambiguous tasks. We introduce Faithful Logic-Aided Reasoning and Exploration (FLARE), which uses LLMs to plan solutions, formalize queries into logic programs, and simulate code execution through multi-hop search without external solvers. Our method achieves SOTA results on $\mathbf{7}$ out of $\mathbf{9}$ diverse reasoning benchmarks and $3$ out of $3$ logic inference benchmarks while enabling measurement of reasoning faithfulness. We demonstrate that model faithfulness correlates with performance and that successful reasoning traces show an $18.1\%$ increase in unique emergent facts, $8.6\%$ higher overlap between code-defined and execution-trace relations, and $3.6\%$ reduction in unused relations.
[…]
FLARE: Faithful Logic-Aided Reasoning and Exploration
Erik Arakelyan†15   Pasquale Minervini23   Pat Verga4   Patrick Lewis4   Isabelle Augenstein1 1University of Copenhagen   2University of Edinburgh 3Miniml.AI   4Cohere   5NVIDIA earakelyan@nvidia.com   augenstein@di.ku.dk   p.minervini@ed.ac.uk {pat, patrick}@cohere.com
@@ 1 Introduction
Complex Reasoning in natural Question Answering (QA) tasks requires exploring a problem space with formalized facts, relations, commonsense knowledge and logical implications. In line with this, LLMs have been enhanced with CoT (Wei et al., 2022) prompting, which supplements the QA process by generating intermediate reasoning chains given a set of in-context examples (Brown et al., 2020a), as shown in fig.˜1. This allowed for advancement in commonsense (Madaan et al., 2022), symbolic (Wang et al., 2022; Sprague et al., 2024) and mathematical (Jie et al., 2023) reasoning. Although CoT allows for a problem exploration in natural language steps, such an approach has been shown to cause performance degradation for reasoning tasks involving multi-step planning (Valmeekam et al., 2022; Suzgun et al., 2023), problem exploration (Yao et al., 2022), and arithmetic tasks (Hendrycks et al., 2021a; Madaan and Yazdanbakhsh, 2022a). These discrepancies arise as CoT suffers from a limited ability to decompose, search, verify and backtrack using intermediate rationale chains (Yao et al., 2022), cascading hallucinations and errors (Ling et al., 2023) and that natural language might not be an optimal representation for describing the reasoning process (Li et al., 2024). Simultaneously, LLM output has been shown to be unfaithful and inconsistent w.r.t. the intermediate CoT rationale (Jacovi et al., 2024; Lanham et al., 2023b; Turpin et al., 2023).
@@
To overcome these problems, we propose Faithful Logic-Aided Reasoning and Exploration (FLARE), an interpretable method that allows for planning, fuzzy reasoning, and traversing the problem space with backtracking, exact task decomposition, and measuring faithfulness. In FLARE, given a natural language query, we prompt an LLM to sequentially generate a plan that includes an analysis and the logical steps necessary for formalising and answering the question, a logic programming (Wielemaker et al., 2012) code that allows formalising the query into a set of facts, relations and their composition forming the space for exploring that query and the search, which is an LLM-generated code execution simulation. An illustration of FLARE can be seen in fig.˜1. This work focuses on models that have not been explicitly trained on CoT on other reasoning traces, as these models have been shown to struggle with generalisation towards differing reasoning paradigms (Chen et al., 2024), consistency in intermediate reasoning steps (Wang et al., 2025) and instruction following (Zhang et al., 2025). In our framework, the generated code must not be consistently executable by an external solver, allowing for the soft-formalisation of natural language. Although we see that even generalist LLMs are able to produce executable code in $\geq 50\%$ of cases. FLARE allows us to measure the faithfulness of the outcome w.r.t. the simulated code execution by directly comparing the search paths produced by the external solver to that LLM generation. This comparison also allows for pinpointing model hallucinations and inconsistencies. We systematically study the effectiveness of our method using $4$ general-purpose LLMs of varying scales across $9$ diverse QA and $3$ logical inference benchmarks, covering Math World Problems, Multi-hop QA, Relation inference, deductive and analytical reasoning and show that our method achieves state-of-the-art results in $7$ out of $9$ QA datasets and $2$ out of $3$ logic datasets in comparison to CoT, F-CoT and Logic-LM. We also show that the method is competitive for models tuned for coding, with an average overall increase of $16\%$ over F-Cot and $10\%$ over CoT. Our key contributions are the following: (i) We introduce FLARE, a novel paradigm for logic-aided and interpretable formalisation and search over the problem space in QA and logic reasoning tasks. (ii) We perform a systematic evaluation across $9$ QA and $3$ logical inference benchmarks and $4$ models of varying scales, showing the advantages of using FLARE for QA in a few-shot setup over prior approaches. (iii) The modularity of FLARE allows defining a simple ingrained method for measuring model faithfulness, which is further shown to be strongly correlated with performance. (iv) We further show that using FLARE allows us to interpretably and rigorously detect hallucinations along with sub-optimal and inconsistent reasoning patterns.
@@ 2 Related Work
Reasoning in Natural Language
Few-shot prompting (Brown et al., 2020b) improves LLM reasoning, and extensions like Chain-of-Thought (CoT)(Wei et al., 2022), “think step by step”(Kojima et al., 2022), and Least-to-Most (Zhou et al., 2023) explicitly decompose problems into intermediate steps. Despite their promise, these methods exhibit arithmetic errors (Lewkowycz et al., 2022; Hendrycks et al., 2021b) and logical inconsistencies (Madaan and Yazdanbakhsh, 2022b). Planning-based variants introduce a separate plan–execute loop (Yao et al., 2023b; Wang et al., 2023a). The plan stage in FLARE draws on these ideas but focuses on generating a natural-language strategy for later formalisation into code.
@@
Recent work augments LLM reasoning by explicitly searching the problem space. Self-consistency decoding (Wang et al., 2023b) samples multiple chains of thought and selects the majority answer, while Tree-of-Thoughts (ToT; Yao et al., 2023a) performs tree-structured exploration with LLM-evaluated states. Later methods adapt classical search—DFS, BFS (Besta et al., 2024), A(Lehnert et al., 2024), and hybrids(Gandhi et al., 2024)—via direct tuning, imitation learning (Yang et al., 2022), or few-shot prompting (Zhang et al., 2024). So far, evaluations focus on toy puzzle and algorithmic domains such as the 24 Game, Countdown, Sorting, mazes, and Sokoban (Yang et al., 2022; Wikipedia, 2024; Besta et al., 2024; Lehnert et al., 2024). Although the search module of FLARE shares this multi-path exploration spirit, it targets more general tasks and yields interpretable multi-hop reasoning via simulated code execution.
@@
Another research direction explores formalising natural language queries into code (Gao et al., 2023; Li et al., 2024) or pseudo-code (Chae et al., 2024; Gandhi et al., 2024). This enables translating queries into strict structures, delegating reasoning and search to deterministic solvers such as Python Chen et al. (2023), PDDL Lyu et al. (2023); Liu et al. (2023), or DataLog Lyu et al. (2023). Models can synthesize programs (Austin et al., 2021; Nijkamp et al., 2023) and benefit from code in numerical and algorithmic reasoning (Chen et al., 2023; Gao et al., 2023), yet their use for general QA remains underexplored. This is due to the challenge of translating natural language into strictly executable code (Wu et al., 2022), the syntactic rigidity of underrepresented programming languages during pre-training (Liu et al., 2024), and the need for models explicitly tuned for coding (Chen et al., 2021). Additionally, relying on external solvers restricts soft reasoning over commonsense knowledge and implications. In FLARE, we formalise queries as logic programs in Prolog during the code generation step but do not require executability or external solvers at inference. This allows LLMs to simulate code execution via soft reasoning over logic-based traversals—similar to Prolog—while avoiding the need for code-specific tuning.
@@
An explanation is considered faithful if it explicitly and accurately describes the reasoning process of the model during inference (Gilpin et al., 2018; Jacovi and Goldberg, 2020). In the context of prompting techniques such as CoT, we are interested in the faithfulness of the intermediate reasoning chains towards the final output. Faithful intermediate reasoning chains should not just look plausible (Herman, 2017) but have exact reflections of the problem exploration and reasoning used to arrive at the final answer. Natural language reasoning chains prevalent in CoT and similar methods are shown to be unfaithful, either masking the reasoning biases (Turpin et al., 2023) of the model or outright ignoring the intermediate reasoning (Lanham et al., 2023a). In FLARE, we introduce a method to seamlessly measure the faithfulness of the final outcome w.r.t. completed search.
@@ 3.1 LLM-Simulated Search
For each query $\mathcal{Q}$, given an LLM $\mathcal{M}$, we initially use instructions $\mathcal{I}^{\mathcal{P}}$ to prompt it to generate a plan $\mathcal{P}$, which should be comprised of task explanation, analysis and a plan for further formalising the query. An example of this can be seen in the plan section in fig.˜1. We use in-context few shot examples $\mathcal{E}_{\mathcal{P}}$ of such plan generations for obtaining the final plan:
$\mathcal{P}_{i}\sim p_{\mathcal{M}}(T^{\mathcal{P}}_{i}\mid T^{\mathcal{P}}_{:i-1},\mathcal{E}_{\mathcal{P}},\mathcal{Q},\mathcal{I}^{\mathcal{P}}),$ (1)
[…]
@@
After generating the plan, we use instructions $\mathcal{I}^{\mathcal{C}}$ to prompt the LLM $\mathcal{M}$ to generate a Prolog code $\mathcal{C}$, an example of which can be seen in fig.˜1. We append executable code generation samples $\mathcal{C}_{\text{sample}}$ to the previous in-context examples $\mathcal{E}_{\mathcal{P}}$ and obtain few-shot code generation demonstrations $\mathcal{E}_{\mathcal{C}}=[\mathcal{E}_{\mathcal{P}};\mathcal{C}_{\text{sample}}]$
$\mathcal{C}_{i}\sim p_{\mathcal{M}}(T^{\mathcal{C}}_{i}\mid T^{\mathcal{C}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}})$ (2)
$\mathcal{F}_{\text{\emph{code}}},\mathcal{R}_{\text{\emph{code}}},\mathcal{G}_{\text{\emph{code}}}=\text{EXTRACT}(\mathcal{C}_{i}),$
[…]
@@
Table 2: Comparison of results across datasets for ChatGPT (gpt-3.5-turbo) and GPT-4 (gpt-4o) using Standard, CoT, Logic-LM, FLARE, and $\text{FLARE}_{SR=2}$ approaches. SR=2 refers to a maximum of 22 iterations of code self-refinement.
@@
After generating the logic-programming code, we want to simulate program execution by generating a problem space traversal trace with our LLM $\mathcal{M}$. We use instructions $\mathcal{I}^{\mathcal{S}}$ and update our in-context samples by appending search traces $\mathcal{S}_{\text{sample}}$ constructed from Prolog execution of sample codes $\mathcal{C}_{\text{sample}}$, i.e. $\mathcal{E}_{\mathcal{S}}=[\mathcal{E}_{\mathcal{C}};\mathcal{S}_{\text{sample}}]$:
$\mathcal{S}_{i}\sim p_{\mathcal{M}}(T^{\mathcal{S}}_{i}\mid T^{\mathcal{S}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}},\mathcal{C},\mathcal{I}^{\mathcal{S}})$ (3)
$\mathcal{A}_{\text{\emph{search}}},\mathcal{F}_{\text{\emph{search}}},\mathcal{R}_{\text{\emph{search}}}=\text{EXTRACT}(\mathcal{S}_{i}),$
@@
$\mathcal{A}_{\text{\emph{Final}}}\sim p_{\mathcal{M}}(T^{\mathcal{A}}_{i}\mid T^{\mathcal{A}}_{:i-1}\mathcal{E}_{\mathcal{C}},\mathcal{Q},$ (4)
$\mathcal{I}^{\mathcal{P}},\mathcal{P},\mathcal{I}^{\mathcal{C}},\mathcal{C},\mathcal{I}^{\mathcal{S}},\mathcal{S},\mathcal{I}^{\mathcal{A}})$
The prompts used for generating each part in FLARE can be seen in section˜A.1 along with a complete example in table˜9 and a pseudo-code in algorithm˜1.